OpenStudy (cherrilyn):

Integrate: dx/(x^2-4)^2

OpenStudy (cherrilyn):

\[\int\limits_{}^{} (dx)/ (x^{2} - 4)^{2}\]

OpenStudy (cherrilyn):

anyone? integration?

OpenStudy (anonymous):

hmm.. i thnk it should be (x-2)^2 and (x+2)^2 then solve it by \[A \div(x-2) + B \div (x-2)^{2}\]

OpenStudy (anonymous):

and so on...

OpenStudy (lgbasallote):

can you use trigonometric substitution? @cherrilyn

OpenStudy (cherrilyn):

yeah im just not sure which trig sub I should use

OpenStudy (lgbasallote):

i mean are you allowed to use

OpenStudy (callisto):

I'm trying to use trigo sub. It turns out to be something... not adorable :(

OpenStudy (cherrilyn):

is wolframalpha's answer correct? I want to know the steps to reach the answer

OpenStudy (lgbasallote):

hmmm yeah not pretty...i'll go with @vedic 's solution \(\LARGE \frac{A}{x+2} + \frac{B}{x-2} + \frac{C}{(x+2)^2} + \frac{D}{(x-2)^2}\) though it's not looking pretty either :/

OpenStudy (anonymous):

@cherrilyn you can click the 'show steps' button

OpenStudy (lgbasallote):

though @psujono you have to admit..wolfram's answers are crazy manipulations :p

OpenStudy (callisto):

If you use trigo. sub. Let x= 2secu dx = 2 secu tanu du \[\int \frac{dx}{(x^2-4)^2} = \int \frac{2 secu tanudu}{((2secu)^2-4)^2}=\frac{1}{8} \int \frac{secu tanudu}{(tan^2u)^2} \]\[=\frac{1}{8} \int \frac{secu du}{(tan^3u)} =\frac{1}{8} \int cscucot^2u du = -\frac{1}{8} \int cotu d(cscu) \]\[= -\frac{1}{8} [cotucscu-\int cscu d(cotu)] =-\frac{1}{8} [cotucscu-\int csc^3u du] \] Then I can't continue :|

OpenStudy (callisto):

Sorry, last step is: \[=-\frac{1}{8} [cotucscu+\int csc^3u d(u)]\]